What is the current in a wire loop of resistance 16 Ω through which the magnetic flux changes from zero to 84 Wb in 3 s?


1
Expert's answer
2021-11-19T10:48:53-0500

Resistance of circuit (R)=16Ω

"Time (t)= 3\times 10^{-3} \;sec"

As we know that, emf is given as-

"e=\frac{dϕ}{dt}=\frac{84-0}{3\times10^{-3}}=28\times10^3\;V"

Therefore,

Current passing through the circuit is-

"I=\frac{e}{R}=\frac{28\times10^3}{16}=1750\;A"


"q=I\times t=1750\times 3\times 10^{-3}=5.250 \;C"


Hence the charge passed through the circuit during the given time is 5.25 C.